# Diffusion Posterior Sampling for Nonlinear CT Reconstruction

### v1

@@ 2.3 Reconstruction using Diffusion Posterior Sampling
For the reverse-time diffusion, we plug in the likelihood score function and generate the samples with the prior score function from the trained score model. The algorithm includes a time-dependent coefficient scheduler $\lambda(t)$, which balances the weighting between the likelihood score function and the prior score function over the process of reverse-time diffusion. Following Chung et al. [13], taking the derivative of the prior score function $\mathbf{s_{\theta^{*}}}(\mathbf{x_{t}},t)$ with respect to the input $\mathbf{x_{t}}$ is accomplished by calculating the Jacobian matrix using backward propagation in the score model. We adopted the standard Variance Preserving Stochastic Stochastic Differential Equations (VP-SDEs) to discretize the reverse process into $N$ bins. The final DPS Nonlinear reconstruction algorithm is summarized in Algorithm 1.
$\mathbf{x_{1}}\sim N(0,I)$,  $\mathbf{\hat{y}}\leftarrow I_{0}\exp(-\mathbf{A}\mathbf{x_{y}})$,  $\mathbf{y}\sim N(\mathbf{\hat{y}},\mathbf{\hat{y}})$,  $\lambda(t)$$\Delta t\leftarrow\frac{1}{N}$;
[…]
        $t\leftarrow(i+1)/N$ $\mathbf{\hat{s}}\leftarrow\mathbf{s_{\theta}^{*}}(\mathbf{x_{t}},t)$ $\mathbf{\hat{x}_{0}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x_{t}}+(1-\bar{\alpha}_{t})\mathbf{\hat{s}})$ $\mathbf{\Delta_{x_{t}}}\leftarrow-\mathbf{A^{T}}\exp(-\mathbf{A}\mathbf{\hat{x}_{0}}){\mathbf{y^{-1}}}I_{0}(\mathbf{y}-I_{0}\exp(-\mathbf{A}\mathbf{\hat{x}_{0}}))$ $\mathbf{\Delta_{x_{0}}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(1+(1-\bar{\alpha}_{i})\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}}))\;\;\;\#\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}})\;is\;taking\;the\;backpropagation_{\theta^{*}}(\mathbf{\hat{s}})\>w.r.t\>\mathbf{x_{t}}$ $\mathbf{dx}\leftarrow[\mathbf{f}(t)\mathbf{x_{t}}+g^{2}(t)(\mathbf{\hat{s}}+\lambda(t)\mathbf{\Delta_{x_{0}}}\mathbf{\Delta_{x_{t}}})]dt$ $\mathbf{z}\sim N(0,I)$ $\mathbf{x_{t-\Delta t}}\leftarrow\mathbf{x_{t}}-\mathbf{dx}+g(t)\sqrt{\Delta t}\mathbf{z}$
end for
@@
$\mathbf{d}\mathbf{x_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)(\nabla_{\mathbf{x_{t}}}\log p(\mathbf{x_{t}})+\nabla_{\mathbf{x_{t}}}\log p_{j}(\mathbf{y_{j}}|\mathbf{x_{t}})]dt+g(t)\mathbf{d}\mathbf{\bar{w}_{t}}\;\;\;t\in(0,T]\;\;\;j\in\{1,2,3,...,N_{os}\},$ (13)
where $j$ is the current index of subset. The subset $j$ can be calculated by $mod(i,N_{os})$, which is taking the remainder of current index of iteration $i$ divided by number of total order subsets. The likelihood term for each order subset $\nabla_{\mathbf{x_{t}}}\log p_{j}(\mathbf{y_{j}}|\mathbf{x_{t}})$ is updated through the following equation
$\nabla_{\mathbf{x_{0}}}\log p_{j}(\mathbf{y_{j}}|\mathbf{x_{0}})=-\mathbf{A_{j}}^{\text{T}}D\{\exp(-\mathbf{A_{j}}\mathbf{x_{0}})\}\mathbf{B}^{\text{T}}D\{1/\mathbf{y_{j}}\}(\mathbf{y_{j}}-\mathbf{B}\exp(-\mathbf{A_{j}}\mathbf{x_{0}}))\;\;\;j\in\{1,2,3,...,N_{os}\},$ (14)
@@
$\mathbf{x_{1}}\sim N(0,I)$,  $\mathbf{\hat{y}_{j}}\leftarrow I_{0}\exp(-\mathbf{A_{j}}\mathbf{x_{y}})$$\mathbf{y_{j}}\sim N(\mathbf{\hat{y}_{j}},\mathbf{\hat{y}_{j}})$,  $\Delta t\leftarrow\frac{1}{N}$$\lambda(t)$$j\in\{1,2,3,...,N_{os}\}$;
[…]
        $t\leftarrow(i+1)/N$ $j\leftarrow mod(i+1,N_{os})$ $\mathbf{\hat{s}}\leftarrow\mathbf{s_{\theta}^{*}}(\mathbf{x_{t}},t)$ $\mathbf{\hat{x}_{0}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x_{t}}+(1-\bar{\alpha}_{t})\mathbf{\hat{s}})$ $\mathbf{\Delta_{x_{t}}}\leftarrow-\mathbf{A_{j}^{T}}\exp(-\mathbf{A_{j}}\mathbf{\hat{x}_{0}})\mathbf{y_{j}^{-1}}I_{0}(\mathbf{y_{j}}-I_{0}\exp(-\mathbf{A_{j}}\mathbf{\hat{x}_{0}}))$ $\mathbf{\Delta_{x_{0}}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(1+(1-\bar{\alpha}_{i})\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}}))\;\;\;\#\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}})\;is\;taking\;the\;backpropagation_{\theta^{*}}(\mathbf{\hat{s}})\>w.r.t\>\mathbf{x_{t}}$ $\mathbf{dx}\leftarrow[\mathbf{f}(t)\mathbf{x_{t}}+g^{2}(t)(\mathbf{\hat{s}}+\lambda(t)\mathbf{\Delta_{x_{0}}}\mathbf{\Delta_{x_{t}}}]dt$ $\mathbf{z}\sim N(0,I)$ $\mathbf{x_{t-\Delta t}}\leftarrow\mathbf{x_{t}}-\mathbf{dx}+g(t)\sqrt{\Delta t}\mathbf{z}$
end for
@@
2.4 Evaluation of DPS Nonlinear and OS-DPS Nonlinear
We evaluate DPS Nonlinear in two different scenarios: 1) Fully sampled low-dose CT reconstruction; and 2) Sparse-view CT reconstruction. For comparison, we compute standard filtered-backprojection (FBP) and model-based iterative reconstructions (MBIR). For the model-based iterative reconstructions, we apply our nonlinear physical model to maximum-likelihood reconstruction using gradient ascent with a fixed step size. For each approach we report Peak Signal-to-Noise ratio (PSNR) and Structured Similarity (SSIM). Additionally, we evaluate OS-DPS Nonlinear in the fully sampled low-dose CT reconstruction scenario and in a limited computation case. For all the experiments, we emulate a system geometry with a source-to-detector distance of 1500 mm, and a source-to-axis distance of 1000 mm. We choose the simple case of $\mathbf{B}=I_{0}$, where $I_{0}$ is a scalar defining the mean number of photons/detector pixel in barebeam. Projections are one-dimensional with 1024 - 0.8 mm pixels. All the systems are implemented using LEAP projectors/backprojectors[26] and PyTorch. Detailed settings and parameters are defined in the following experiments.
@@ 2.4.1 Low-dose CT Reconstruction Comparisons
For the low-dose CT reconstruction scenario, $I_{0}=10^{3}$ photons/pixel with 360 projections covering $360^{\circ}$. The lambda scheduler was set to $\lambda(t)=10^{-3.1t-0.7}$ where the $t$ is the reverse timestep from 1 to 0. The number of iterations in the reverse-diffusion is set to $1000$. For the MBIR, we adopt $40000$ gradient ascent iterations with step size equals to $6\times 10^{-7}$.
@@ 2.4.2 Sparse view CT reconstruction Comparisons
Sparse view CT reconstruction is applied to a system with $I_{0}=10^{5}$ photons/pixel with 30 projects cover $360^{\circ}$. In this case, the lambda scheduler is set to $\lambda(t)=10^{-4.3t-1.5}$ and the number of iterations in the reverse diffusion process is set to $1000$. For MBIR, we adopt $150000$ gradient ascent iterations with step size equals to $3\times 10^{-8}$.
@@ 2.4.3 OS-DPS Nonlinear Comparisons
For evaluation of OS-DPS Nonlinear, use the same scenario as low-dose CT reconstruction. The system is defined by $I_{0}=10^{3}$ photons/pixel with 360 projections covering $360^{\circ}$. We consider 1, 6, 12, 24, 36, 48, and 60 ordered subsets (OS) for the comparison. The lambda scheduler remains the same as the low-dose CT reconstruction when the OS = 1 and is modified to $10^{-5.5t+1.7}$ for OS = 6, $10^{-6.6t+2.8}$ for OS = 12, $10^{-7.5t+3.7}$ for OS = 24, $10^{-8.4t+4.6}$ for OS = 36, $10^{-8.8t+4}$ for OS = 48, and $10^{-9.2t+5.4}$ for OS = 60. According to the Equation (6), $\lambda(t)$ should not be greater than 1 during the whole reverse-time process. Therefore, any lambda in the reverse process that is greater than 1 is thresholded back to 1. We first apply the same $1000$ iterations in the reverse-diffusion to evaluate differences. In addition to PSNR and SSIM, we proposed another way of comparing the results generated by diffusion posterior sampling. For a single input image, we run the posterior sampling to generate 32 results. Based on these 32 generated results, we take the mean and standard deviation of them to get the sample mean image and standard deviation image. We further subtract the mean from the ground truth image to get a bias image. We compare sample mean image, bias image, and standard deviation images between different numbers of ordered subsets.
To further test our method, we apply our OS-DPS Nonlinear under a different scenario where the system is defined by $I_{0}=10^{4}$ photons/pixel with $200$ projections over $360^{\circ}$. To consider the potential of ordered subsets under limited computation time, we compare results with a relatively low fixed number of $200$ iterations/time steps. The lambda scheduler is set to $10^{-3.4t-1.4}$ for OS = 4, $10^{-3.7t+1.1}$ for OS = 8, and $10^{-4.1t-0.7}$ for OS = 20.
@@ 3.1 Low-dose CT Reconstruction
Figure 1 shows reconstructions in the low-dose CT scenario. We see that the proposed DPS reconstruction yields a sharp image with recognizable anatomical structures, indicating that the proposed method is capable of guiding the unconditional diffusion generation through the posterior sampling. In contrast, the FBP reconstruction has heavy streaking associated with the very low x-ray fluence (particularly along the long axis and the spine where photon starvation effects are more pronounced.) The MBIR approach is both noisy and with residual blur, obscuring many fine details. We found the PSNR of FBP was $27.67$ dB, MBIR was $32.26$ dB, and the DPS approach was $54.55$ dB. For the SSIM, DPS achieved $0.7224$ while FBP was $0.0742$ and MBIR was $0.2082$. These quantitative measures exhibit the same rank ordering of our subjective image quality assessment.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Low-dose CT reconstructions comparing FBP, MBIR, and the proposed DPS Nonlinear approach. All images share a common colormap with W/L = 2000/0 HU.
Looking closely at the DPS reconstruction for low-dose CT reconstruction, we note that the generative aspect of the DPS approach is evident. There are features in the DPS reconstruction that do not match exactly those found in the ground truth. Such hallucinations are common with these generative approaches - particularly when applied to very low-fidelity data as in this scenario.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Additional DPS sample reconstructions from the same data illustrating variability in the outputs.
Figure 2 further explores some of the generative aspects of the proposed DPS approach. Among several sample outputs applied to the same data, one can see both common features and variable features in the reconstruction. In general, DPS Nonlinear is more stable in the generation of low frequency general structures than the high frequency details. In particular, the fine details in the lung tend to be most variable with more consistency in major high contrast structures. This is particularly evident from a variance image computed from 32 of these DPS examples. Note the increased variability at tissue boundaries and in the smaller features in the lungs.

### v2

@@ 2.3 Reconstruction using Diffusion Posterior Sampling
For the reverse-time diffusion, we plug in the likelihood score function and generate the samples with the prior score function from the trained score model. The algorithm includes a time-dependent coefficient scheduler $\lambda_{t}$, which balances the weighting between the likelihood score function and the prior score function over the process of reverse-time diffusion. We designed a heuristic weighting scheduler $\lambda_{t}$ that can greatly stabilize the reverse process in our experiments:
$\lambda_{t}=\frac{k}{||\nabla_{\mathbf{x_{t}}}\log p(\mathbf{y}|\mathbf{x_{t}})||^{2}_{2}},$ (10)
where weighting coefficient $k$ is a scalar that is weighted by the magnitude of the gradient likelihood and that can be tuned manually under different scenarios. Following Chung et al. [13], taking the derivative of the prior score function $\mathbf{s_{\theta^{*}}}(\mathbf{x_{t}},t)$ with respect to the input $\mathbf{x_{t}}$ is accomplished by calculating the Jacobian matrix using backward propagation in the score model. We adopted the standard Variance Preserving Stochastic Stochastic Differential Equations (VP-SDEs) to discretize the reverse process into $N$ bins. The final DPS Nonlinear reconstruction algorithm is summarized in the following Algorithm 1.
1:$\mathbf{x_{1}}\sim\mathcal{N}(0,I)$,  $\mathbf{\hat{y}}\leftarrow I_{0}\exp(-\mathbf{A}\mathbf{x_{y}})$,  $\mathbf{y}\sim Poisson(\mathbf{\hat{y}})$,  $\lambda(t)$$\Delta t\leftarrow\frac{1}{N}$
2:# Diffusion Posterior Sampling
3:for $i=N-1$ to 00 do:
4:     $t\leftarrow(i+1)/N$
5:     $\mathbf{\hat{s}}\leftarrow\mathbf{s_{\theta}^{*}}(\mathbf{x_{t}},t)$
6:     $\mathbf{\hat{x}_{0}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x_{t}}+(1-\bar{\alpha}_{t})\mathbf{\hat{s}})$
7:     $\mathbf{\Delta_{x_{t}}}\leftarrow-\mathbf{A}^{T}\mathbf{y}+\mathbf{A}^{T}D\{I_{0}\exp(-\mathbf{A}\mathbf{{x}_{0}})\}$
8:     $\mathbf{\Delta_{x_{0}}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(1+(1-\bar{\alpha}_{t})\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}}))$
9:     # $\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}})$ is taking the backpropagation w.r.t $\mathbf{x_{t}}$
10:     $\mathbf{dx}\leftarrow[\mathbf{f}(t)\mathbf{x_{t}}+g^{2}(t)(\mathbf{\hat{s}}+\lambda(t)\mathbf{\Delta_{x_{0}}}\mathbf{\Delta_{x_{t}}})]dt$
11:     $\mathbf{z}\sim\mathcal{N}(0,I)$
12:     $\mathbf{x_{t-\Delta t}}\leftarrow\mathbf{x_{t}}-\mathbf{dx}+g(t)\sqrt{\Delta t}\mathbf{z}$
13:end for
@@
$\mathbf{d}\mathbf{x_{t}}=[\mathbf{f}(\mathbf{x_{t}},t)+g^{2}(t)(\nabla_{\mathbf{x_{t}}}\mathrm{log}p(\mathbf{x_{t}})+\nabla_{\mathbf{x_{t}}}\mathrm{log}p_{j}(\mathbf{y_{j}}|\mathbf{x_{t}})]dt+g(t)\mathbf{d}\mathbf{\bar{w}_{t}}\;\;\;t\in(0,T]$ (11)
where $j\in\{1,2,3,...,N_{os}\}$ is the current index of subset, $t\in(0,T]$ is the timesteps. The subset $j$ can be calculated by $mod(i,N_{os})$, which is taking the remainder of current index of iteration $i$ divided by number of total order subsets. The likelihood term for each order subset $\nabla_{\mathbf{x_{t}}}\mathrm{log}p_{j}(\mathbf{y_{j}}|\mathbf{x_{t}})$ is updated through the following equation
${\nabla_{\mathbf{x_{0}}}\mathrm{log}p_{j}(\mathbf{y_{j}}|\mathbf{x_{0}})=-\mathbf{A_{j}}^{T}\mathbf{y_{j}}+\mathbf{A_{j}}^{T}\mathbf{D}\{I_{0}\exp(-\mathbf{A_{j}}\mathbf{{x}_{0}})\}.}$ (12)
@@
1:$\mathbf{x_{1}}\sim\mathcal{N}(0,I)$, $\Delta t\leftarrow\frac{1}{N}$, $\lambda(t)$
2:$\mathbf{\hat{y}_{j}}\leftarrow I_{0}\exp(-\mathbf{A_{j}}\mathbf{x_{y}})$$\mathbf{y_{j}}\sim Poisson(\mathbf{\hat{y}_{j}})$$j\in\{1,2,3,...,N_{os}\}$
3:for $i=N-1\ \text{to}\ 0$ do
4:     $t\leftarrow(i+1)/N$
5:     $j\leftarrow mod(i+1,N_{os})$
6:     $\mathbf{\hat{s}}\leftarrow\mathbf{s_{\theta}^{*}}(\mathbf{x_{t}},t)$
7:     $\mathbf{\hat{x}_{0}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x_{t}}+(1-\bar{\alpha}_{t})\mathbf{\hat{s}})$
8:     $\mathbf{\Delta_{x_{t}}}\leftarrow-\mathbf{A_{j}}^{T}\mathbf{y_{j}}+\mathbf{A_{j}}^{T}\mathbf{D}\{I_{0}\exp(-\mathbf{A_{j}}\mathbf{{x}_{0}})\}$
9:     $\mathbf{\Delta_{x_{0}}}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(1+(1-\bar{\alpha}_{t})\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}}))$
10:     # $\nabla_{\mathbf{x_{t}}}\mathbf{\hat{s}}(\mathbf{x_{t}})$ is taking the backpropagation w.r.t $\mathbf{x_{t}}$
11:     $\mathbf{dx}\leftarrow[\mathbf{f}(t)\mathbf{x_{t}}+g^{2}(t)(\mathbf{\hat{s}}+\lambda(t)\mathbf{\Delta_{x_{0}}}\mathbf{\Delta_{x_{t}}}]dt$
12:     $\mathbf{z}\sim\mathcal{N}(0,I)$
13:     $\mathbf{x_{t-\Delta t}}\leftarrow\mathbf{x_{t}}-\mathbf{dx}+g(t)\sqrt{\Delta t}\mathbf{z}$
14:end for
2.4 Evaluation Studies
We evaluate DPS Nonlinear in three scenarios: 1) Fully sampled low-dose CT reconstruction; 2) Sparse-view CT reconstruction; and 3) CT reconstruction across varying noise levels. These three scenarios share the same CT system geometry with a source-to-detector distance of 1500 mm, and a source-to-axis distance of 800 mm. Projections are one-dimensional with 1024 - 1.556 mm pixels. For the fully sampled low-dose CT reconstruction, $I_{0}=10^{3}$ photons/pixel with 360 projections covering $360^{\circ}$. For the Sparse view CT reconstruction, $I_{0}=10^{5}$ photons/pixel with 30 projections cover $360^{\circ}$. For CT reconstruction across multiple noise levels. We consider the fully sampled CT systems with varied incident fluence, where $I_{0}$ equals to $10^{2}$, $10^{3}$, $10^{4}$, and $10^{5}$ photons/pixel with 360 projections covering $360^{\circ}$.
For the fully sampled low-dose and sparse-view CT reconstructions, we compute traditional filtered-backprojection (FBP) and model-based iterative reconstructions (MBIR) (with and without total variation regularization [26]). Additionally, we compare DPS Nonlinear to two related diffusion approaches to investigate potential advantages. In particular, we seek to study the impact of the unsupervised model training and the explicitly nonlinear model of the proposed approach. Toward that end we select: 1) A diffusion technique that uses supervised training and conditioning on measurement data - DOLCE [23] and, 2) Diffusion posterior sampling based on a linearized measurement model [13] (which we will refer to as DPS Linear). The linearized forward measurement is defined as follows:
$\mathbf{l}=\mathbf{A}\mathbf{x},\;\;\;\text{where}\;\;\;\mathbf{l}=-\log(\mathbf{y}/I_{0}),$ (13)
where DPS Linear aims to solve the inverse of the linear reconstruction problem using diffusion posterior sampling. We further report the inference runtime of all three diffusion methods in the fully sampled low-dose scenario Additional evaluations of OS-DPS Nonlinear were conducted using the fully sampled low-dose CT reconstruction scenario in a limited computation case.
For each approach we report Peak Signal-to-Noise ratio (PSNR) and Structured Similarity (SSIM). For the evaluation of diffusion posterior sampling based methods like DPS Linear and DPS Nonlinear, we propose an additional generation evaluation method: For the same noisy measurement, we run the posterior sampling to generate multiple results and calculate the bias image $\mathbf{x}_{bias}$ and standard deviation image $\mathbf{x}_{std}$ defined as follows:
$\mathbf{x}_{bias}=|\mathbb{E}[\mathbf{\hat{x}}]-\mathbf{x_{0}}|$ (14)
$\mathbf{x}_{std}=\sqrt{\mathbb{E}[(\mathbf{\hat{x}}-\mathbb{E}[\mathbf{\hat{x}}])^{2}]}$ (15)
where $\hat{x}$ is all the generated sample images, $\mathbf{x_{0}}$ is the ground truth. We can summarize the bias and standard deviation by taking the root mean square of bias image $\mathbf{x}_{bias}$ and mean of standard deviation image $\mathbf{x}_{std}$, respectively. All the methods and systems are implemented using PyTorch and LEAP projectors / backprojectors[27] on the platform of NVIDIA GeForce RTX 3090Ti GPU with Intel(R) Xeon(R) CPU E5-2697 v4. Detailed settings are defined in the following experiments.
@@ 2.4.1 Low-dose CT Reconstruction Comparisons
For MBIR, we adopt $10000$ iterations using Adam gradient optimizer with initial step size set to $0.01$ with total variance regularization (TV). This large number of iterations and relatively small step size was used to ensure a highly converged image estimate. The TV weighting was selected as $6\times 10^{-6}$ based on a subjective image quality assessment. DOLCE is a supervised diffusion-based method which requires a FBP or MBIR volume as an input in addition to projection data. We implemented DOLCE using the same training and validation sets as DPS Nonlinear with a fully sampled low-dose FBP as an additional input. A low-dose FBP was also used as a conditional input for subsequent evaluations. For the DPS Linear diffusion model, we adopted the same model weights as DPS Nonlinear since only the conditional reverse process is different. A standard DDPM sampler [9] with $1000$ reverse sampling steps was applied for all three diffusion-based methods (DPS Nonlinear, DPS Linear, DOLCE). For fair comparison between DPS Linear and DPS Nonlinear, we evaluated 16 different weighting coefficients $k$ in equation (10) from $70$ to $490$ with an interval of 30, and computed the root mean squared bias and standard deviation. (Based on the comparison of bias and standard deviation, we selected $k=310$ as the best weighting for image comparison.)
@@ 2.4.2 Sparse view CT reconstruction Comparisons
We apply MBIR to the low-dose scenario with same parameters as the fully sampled scenario, except that the TV was chosen empirically to be $2\times 10^{-5}$. To evaluate the ability of a supervised diffusion approach to operate in a different imaging scenario than it was trained, we evaluate the performance of DOLCE (trained on the fully sampled scenario) applied to the sparse-view scenario. Sparse-view FBP is used as a conditional input for DOLCE. Again, for DPS Linear and DPS Nonlinear, different $k$ coefficients were evaluated from 200 to 1320 with an interval of 80. (We found $k=840$ based on the consideration of both root mean squared bias and standard deviation value.)
2.4.3 Reconstruction across a range of fluence/noise levels
We evaluate DPS Nonlinear across a range of fluence levels. We optimize the weighting coefficient individually for each fluence level and find optimal values at $k=310$ for $I_{0}=10^{2}$, $k=780$ for $I_{0}=10^{3}$, $k=1080$ for $I_{0}=10^{4}$, and $k=1280$ for $I_{0}=10^{5}$ based on RMSB. We generate $400$ DPS samples and showcase the one closest to the mean PSNR (over those samples). We also compute the error map for that sample, and a standard deviation map showing variability over all samples. To further illustrate variability and susceptibility to hallucination as a function of fluence level, we present the variation in both global PSNR (across the entire image) and local PSNR (within a lung region) as a function of fluence. Moreover, we compute error images for the best and worst case samples (based on local PSNR).
@@ 2.4.4 OS-DPS Nonlinear Comparisons
For evaluation of OS-DPS Nonlinear, we consider 1, 3, 6, 12, 18, 24, 36, 48, 60, 72, 120, and 180 ordered subsets (OS) in the fully sampled low-dose scenario for the comparison. The weighting coefficient is set to $k=150$ according to the same optimization strategy as in fully low-dose and sparse-view scenarios. In addition to PSNR and SSIM, we also plot out the bias and standard deviation images across all the ordered subsets.
@@ 3.1 Low-dose CT Reconstruction
Figure 2 shows the image results from the different reconstruction methods for the low dose CT scenario. The three diffusion-based methods yield sharp images with clear anatomical structures. In contrast, the FBP reconstruction has heavy streaking associated with the very low x-ray fluence (particularly along the long axis and the spine where photon starvation effects are more pronounced.) The MBIR approach without regularization is very noisy while total variance regularized version has reduced noise, blurring obscures many of the fine details. The PSNR of our proposed technique was $28.32$ dB, which is $0.67$ dB ($\sim 2\%$) higher than DPS Linear and $0.38$ dB ($\sim 1\%$) higher than DOLCE. For SSIM, DOLCE achieves the best with $0.8529$. Our proposed DPS Nonlinear achieved $0.8350$, which is $0.0179$ lower than DOLCE ($\sim 2\%$), but higher than DPS Linear. While one might expect DOLCE to outperform the other DPS techniques because of the supervised training where system-specific ground truth and degraded images are provided, we see that the relative performance among these diffusion methods is relatively close. We note that all diffusion methods show a degree of false feature generation (particularly in the fine detail structure of the lungs).
Figure 2: Low-dose CT reconstructions comparing FBP, MBIR (without and with TV regularization), DPS Linear, DOLCE, and the proposed DPS Nonlinear approach. All images share a common colormap with W/L = 2000/0 HU.
Figure 3 shows the evaluation of DPS Linear and DPS Nonlinear under different weighting coefficients, $k$ values, in the low-dose scenario. We can see that DPS Nonlinear and DPS Linear share a similar performance curve line for root-mean-squared bias and standard deviation. For the best weighting coefficient, $k=310$, the proposed DPS Nonlinear achieves $\sim 4.7\%$ lower bias compared to DPS Linear while maintaining the similar standard deviation. Such biases have been noted in linearized MBIR reconstructions at low exposures due to the approximations to the Poisson model, and that same kind of bias appears for DPS Linear versus DPS Nonlinear (where no approximation is made).
Figure 3: Bias and standard deviation comparison for low-dose CT reconstructions comparing the proposed DPS Nonlinear and DPS Linear approach. Root-mean-squared bias and standard deviation are reported in Hounsfield units (HU) based on 16 generated samples from diffusion posterior sampling for each weighting coefficient $k$ value. Error bars (±$\pm~{}1.0$ standard deviation) are shown based on the variability over samples. The best weighting coefficient $k$ was $310$ based on the lowest root-mean-squared bias as well as the lowest standard deviation.
Figure 4 further explores some of the generative aspects of the proposed DPS approach. Among several sample outputs applied to the same data, one can see both common features and variable features in the reconstruction. In general, DPS Nonlinear is more stable in the generation of low frequency general structures than the high frequency details. In particular, the fine details in the lung tend to be most variable with more consistency in major high contrast structures. This is particularly evident from a variance image computed from 16 of these DPS examples. Note the increased variability at tissue boundaries and in the smaller features in the lungs.
Figure 4: Additional DPS sample reconstructions and standard deviation image from the same noisy measurement illustrating variability in the outputs. Colormap: W/L = 2000/0 HU for sample image and W/L = 500/250 HU for standard deviation image.